# Towards Computation- and Communication-efficient Computational Pathology

### v1

@@ Results
Figure 4: Extensibility of MAGA to the histopathology-specific encoder. MAGA can be applied to a histopathology-specific encoder or pathology vision foundation model to extract fine-grained domain-specific features from low magnification inputs. Receiver operating characteristic (ROC) curves demonstrate the performance of CLAM (a, d, g), TransMIL (b, e, h) and MAGA-GLTrans (c, f, i) on TCGA-NSCLC using MAGA trained with domain-unspecific (ImageNet pre-trained ResNet50) and domain-specific models (TCGA pre-trained CTransPath), respectively. From top to bottom: ROC curves at 20×20$\times$20 ×, 5×5$\times$5 × without MAGA and 5×5$\times$5 × with MAGA.
MAGA can be extended to the histopathology-specific encoder
@@
Histopathology-specific encoders or foundation models [24, 23] have been proven to be much more effective for WSI analysis in various CPath tasks compared with the feature extractor pre-trained on natural images. MAGA, as a self-supervised model that can align the learned knowledge and restore the information loss between two magnifications, is capable of being extended to the histopathology-specific encoder, making it possible for low magnification inputs. Here, we investigate the extensibility of MAGA on the earliest histopathology-specific model, CTransPath [24] as well as the recently proposed pathology vision foundation model, UNI [23]. We conduct this experiment on the NSCLC subtyping task (TCGA-NSCLC). CLAM, TransMIL and MAGA-Trans are trained at the high magnification 20×20$\times$20 × and the low magnification 5×5$\times$5 × (with and without MAGA).
@@
As we expected, the histopathology-specific pre-trained feature extractors at 20×20$\times$20 × brings appropriate improvements for all three models (Figure 4a-c) compared with the ImageNet pre-trained feature extractor with respect to AUC (CTransPath: CLAM=0.9631, $+7.01\%$, TransMIL=0.9536, $+3.52\%$, MAGA-GLTrans=0.9659, $+0.83\%$) and (UNI: CLAM=0.9790, $+8.06\%$, TransMIL=0.9690, $+5.06\%$, MAGA-GLTrans=0.9673, $+0.97\%$). When reducing the magnification level to 5×5$\times$5 × without the support of MAGA, the classification performance (Figure 4d-f) with three different encoders of the three WSI analysis models decreases due to the information loss. Even though CTransPath without MAGA also faces an obvious performance degradation at 5×5$\times$5 ×, applying a histopathology-specific feature extractor is still a better solution compared with the ImageNet pre-trained one. When training MAGA on three different encoders at the 5×5$\times$5 × magnification level, the performance of three WSI analysis architectures returns to a favorable level, even outperform the model trained at 20×20$\times$20 × with the ImageNet pre-trained feature extractor. The model-agnostic nature of MAGA is a highly significant feature that enables it to be seamlessly integrated with any top-tier, histopathology-specific pre-trained model, resulting in minimal information loss. This capability substantially enhances computational efficiency and communication efficiency, all without compromising the model’s classification performance.
@@
Figure 5: Visualization of the features extracted by MAGA (TCGA-NSCLC). a, The prediction results (top) of one fold on the subtyping task (TCGA-NSCLC, $n=106$) under different magnification levels. Patients are sorted by the ground truth labels. Heatmaps (bottom) demonstrate the similarity of the features between high magnification level and low magnification level (with and without MAGA). We show the feature similarity in slide level, which is the mean of all the patches. Red represents higher feature similarity. b, A sample case WSI (left) of LUSC with pathologist-annotated carcinoma region (yellow border), visualization of the unsupervised clustering of the patch-level features (top right) and the heatmap (bottom right) at 20×20$\times$20 ×, 5×5$\times$5 × with MAGA and 5×5$\times$5 × without MAGA. c, T-SNE visualization of the image embeddings from three different models in the Kather colon dataset [39] with nine tissue types.
MAGA features are reliable and interpretable
@@
Model interpretability and reliability are the most important natures when applying AI models to real-world clinical scenarios. As a feature extractor given low magnification inputs, MAGA faces a significant challenge of information loss, potentially undermining the model’s reliability and interpretability. In this experiment, we investigate the feature quality with and without MAGA, by visualizing them at cohort, slide and patch levels. The visualization of NSCLC subtyping (TCGA-NSCLC) and breast cancer lymph-node metastasis detection (CAMELYON16) are shown in Figure 5 and Extended Figure 4 respectively.
@@
To evaluate and visualize whether MAGA is able to restore the information loss at the low magnification, we quantify the similarity of the features between low magnification 5×5$\times$5 × (with and without MAGA) and high magnification 20×20$\times$20 ×. Specifically, we calculate a slide-level similarity score by averaging the similarity scores (normalized Euclidean distances) across all patches. The heatmaps depicted in Figure 5a indicate that the features extracted from the low magnification images using MAGA are significantly more similar to those from the high magnification images than the features extracted at low magnification without MAGA. This finding suggests that MAGA can refine the features extracted from low magnification images, effectively bridging a substantial information gap of up to 16 times. With those more reliable features, the rate of misclassify has noticeably decreased at 5×5$\times$5 × magnification using MAGA compared with the ones in the 5×5$\times$5 × without MAGA. We also observe that the misclassified samples at 5×5$\times$5 × without MAGA tend to exhibit lower feature similarities, typically falling below the threshold of 0.5. For the CAMELYON16 dataset with fewer training samples and larger color variations, unreliable features can severely impair the prediction model, resulting in significantly degraded classification performance with more misclassified samples (Extended Figure 4).
@@
To further visualize the robustness and the representativeness of MAGA features, and to evaluate whether those features deliver enough semantic information, we present a slide-level visualization in Figure 5b. We first apply an unsupervised clustering approach (K-means) on the patch-level features and group them into four clusters. The cluster maps shown in the top-right row demonstrate that the features extracted from high magnification patches are capable of capturing and reflecting the morphological characteristics accurately with a clear boundary between tumor and normal tissues and a high agreement with the boundary drawn by the pathologist. When reducing the magnification to 5×5$\times$5 ×, the tissue boundaries become unclear and the clustering map is much noisier than the high magnification one. This observation indicates that reducing image resolution can greatly harm the semantic information, resulting in unstable prediction results. By bridging information gap using MAGA, the clustering map of 5×5$\times$5 × with MAGA maintains a high visual consistency with the one of 20×20$\times$20 ×. Next, we plot the heatmaps of the WSI analysis model GLTrans in the top-right row in Figure 5b. The color intensity of the heatmaps correlates with the confidence of the classification outcome for that particular patch. When without MAGA, the lower magnification features lead to less effective heatmaps, indicating that the feature representation of the predictive models is impaired by these unstable and less detailed features. Applying MAGA significantly enhances the effectiveness of features, resulting in heatmaps that are much more consistent with those at higher magnification. The feature space of the image embeddings (Figure 5c) also demonstrates the effectiveness of MAGA.
@@
Figure 6: MAGA-GLTrans on telepathology for intraoperative frozen section diagnosis. a, Local hospital transfers low magnification WSIs to the cloud server and obtains a diagnostic result and an interpretable heatmap in a only few seconds. b-f, Statistical results (b) and the box plots at different magnification levels with respect to the pixel quantity (c), data transfer time (d), model inference time (e) and the diagnostic performance of AUC (f). h, A sample case and its corresponding heatmap as well as the patches with high confidence (red border) and low confidence (blue border).
@@
We evaluate the potential of MAGA-Trans for computer-assisted telepathology in intraoperative frozen section diagnosis, a critical need for hospitals lacking on-site expert pathologists, with a focus on classification performance, computational efficiency, and communication efficiency. This investigation is conducted on the NSCLC subtyping task using the TCGA-NSCLC dataset (frozen sections), which includes 1067 frozen section WSIs of LUAD and 1100 WSIs of LUSC. A 10-fold cross-validation approach is employed, with all patients randomly partitioned into a training set (80%), a validation set (10%), and a test set (10%). To prevent data leakage, all slides from a single patient are assigned to the same set. Classification performance is assessed using the AUC metric. For each WSI, we store each magnification level as a separate file to accurately measure storage consumption, data transfer time, and model inference time across different magnifications. Transmission performance is evaluated under a 200 Mbps bandwidth connection during typical working hours (8:00-18:00) to simulate real-world clinical conditions.
@@
For local hospitals lacking on-site expert pathologists, particularly in underdeveloped regions or countries, we propose an efficient computer-assisted telepathology (CATP) framework tailored for frozen section diagnosis (Figure 6a). This framework enables local hospitals to transmit low-magnification whole slide images (WSIs) to cloud servers and receive computer-assisted diagnostic results alongside interpretable heatmaps within seconds. To rigorously assess the framework’s efficacy, we conducted an ablation study, as depicted in Figure 6b. The results reveal that, as anticipated, higher magnification WSIs require significantly more storage due to their substantially larger pixel counts (20$\times$: 0.503 Gigapixels/WSI vs. 5$\times$: 0.033 Gigapixels/WSI), leading to increased data transfer and model inference times. On average, processing a single WSI takes approximately one minute (59.13 seconds) for file transmission and half a minute (32.32 seconds) for model inference. In real-world clinical settings, these times can vary considerably due to factors such as network bandwidth, latency fluctuations, computational resource availability, concurrent user load, and other environmental variables. Reducing magnification yields remarkable efficiency gains, improving processing speed by more than tenfold. Specifically, data transfer time drops dramatically from 59.13 seconds per WSI at 20$\times$×magnification to just 2.79 seconds at 5$\times$, while model inference time decreases from 32.32 seconds to around 5 seconds per WSI. However, this optimization comes at a significant cost to diagnostic accuracy, as indicated by the sharp decline in AUC values from 0.8995 at 20$\times$ to 0.6267 at 5$\times$, severely limiting clinical utility. Our proposed MAGA-GLTrans framework effectively addresses this trade-off by maintaining high predictive performance (AUC: 0.8418) with only marginal degradation compared to 20$\times$ magnification, while simultaneously mitigating the drawbacks of magnification reduction. By significantly reducing both data transfer and model inference times, this innovative CATP framework greatly enhances the clinical feasibility of intraoperative frozen section diagnosis.
@@ Discussion
We introduce MAGA-GLTrans, a general-purpose WSI analysis architecture, designed for a wide range of computational pathology applications, including cancer subtyping, metastasis detection, and mutation prediction. Beyond its exceptional classification capabilities, MAGA-GLTrans stands out for its computational efficiency, storage-friendly design, extensibility, and communication efficiency. These attributes, whose importance is first emphasized in this study, are often overlooked in the field of computational pathology (CPath). The primary obstacle to advancing CPath towards greater efficiency lies in the fact that current WSI analysis architectures typically process WSIs at high magnifications [40]. This approach is constrained by the limitations of existing computer hardware. While directly reducing image resolution can improve efficiency, it often comes at the cost of performance due to information loss. MAGA addresses this challenge by bridging the information gap and extracting fine-grained features from low-magnification images, thereby achieving an optimal balance between performance and efficiency. Through 10-fold cross-validation experiments on five CPath tasks, our model demonstrates robust classification performance at 5$\times$×magnification, comparable to two state-of-the-art WSI analysis architectures operating at 20$\times$, while reducing computational time by up to 10.7 times, storage overhead by up to 28.6 times, and communication time by up to 21.5 times. We believe this represents a significant and milestone finding that brings CPath technology one step closer to real-world clinical adoption. For instance, a recent study highlights that reviewing WSIs results in a median overall 19% decrease in efficiency per case compared to glass slides [41]. This inefficiency is further exacerbated when pathologists must wait tens of seconds for a decision on a single slide from CPath systems. Such delays are particularly pronounced when analyzing multiple hematoxylin and eosin (H&E) stained diagnostic slides and immunohistochemistry slides, a common scenario for patients with larger tumors. Even if WSIs are processed offline before pathologist review, hospitals with high patient volumes would require substantial computational resources. MAGA-GLTrans addresses this issue by reducing computation turnaround time to just a few seconds, making the computational process virtually seamless for pathologists.
@@
In addition to enhancing local pathology diagnosis, we also investigate the potential of MAGA-GLTrans in enabling telepathology (TP), a field of pathology that utilizes telecommunication technologies to facilitate the transfer of digital pathology slides across geographically distant locations [42]. Intraoperative frozen section diagnosis stands out as one of the most promising applications of TP, particularly in settings where pathologists are scarce, such as hospitals without on-site pathology expertise [43, 44, 45]. Integrating computational pathology techniques to enable computer-assisted telepathology (CATP) for intraoperative diagnosis has the potential to improve diagnostic accuracy [46]. However, the unpredictable data transfer times under complex network conditions and the prolonged inference times associated with processing gigapixel WSIs pose significant challenges, potentially increasing turnaround times and compromising diagnostic efficiency [25]. These issues highlight the need for further validation of CATP’s feasibility. Our evaluation on the frozen section diagnosis dataset (TCGA-NSCLC (frozen sections)) demonstrates that MAGA-GLTrans can significantly reduce the turnaround time required for both data transfer and computation.
@@
Due to the absence of domain-specific encoders before the emergence of foundation models, earlier WSI analysis pipelines typically relied on ImageNet pre-trained models to extract patch-level features as a compromise solution. However, with the accumulation of large-scale histopathology datasets and advancements in computational resources, self-supervised learning on histopathology datasets [24] and the development of visual foundation models [13, 23, 47] have led to significant progress in computational pathology. Notably, histopathology-specific encoders such as CTransPath [24], UNI [23], PLIP [13], Prov-GigaPath [11], MUSK [48] and Virchow [49] have demonstrated superior effectiveness across various CPath tasks compared to encoders trained on natural images. To preserve fine-grained information, most existing histopathology-specific encoders are still trained on high-magnification images. MAGA, however, can be extensible to any histopathology-specific encoder in a self-supervised manner, enabling the extraction of fine-grained features even from low-magnification inputs. Experimental results indicate that MAGA, when trained with a histopathology-specific encoder, further enhances classification performance compared to MAGA trained with ImageNet. Additionally, due to its plug-and-play nature, MAGA can be seamlessly integrated into other WSI analysis architectures, inheriting their strengths while improving their efficiency. For example, when MAGA is applied as the feature extractor for CLAM, the performance remain largely unaffected even at low magnification. This exceptional extensibility of MAGA significantly extends its lifecycle, particularly in light of the rapid pace of algorithmic advancements and iterations in the field.
@@
While the MAGA-GLTrans model has demonstrated remarkable efficiency and accuracy in handling low-magnification inputs, several areas warrant further exploration to unlock its full potential. Firstly, we hypothesize that the model’s capabilities could extend beyond its current resolution limits. By applying the model to even lower magnifications, such as 2.5$\times$, it may be possible to achieve further improvements in computational speed and communication efficiency. However, this approach requires a careful balance between reducing image resolution and preserving essential diagnostic details. Whether this trade-off will compromise the model’s effectiveness at lower magnifications remains an open question that merits thorough investigation. Additionally, due to the lack of large-scale datasets, MAGA was trained on a limited number of WSIs as a compromise. Investigating its performance on a larger-scale dataset would be highly valuable, as it could further enhance the model’s generalization capabilities and robustness. This exploration could provide deeper insights into the scalability and effectiveness of MAGA when applied to more diverse and extensive histopathology data, potentially unlocking even greater performance improvements and broader applicability in computational pathology.
@@
In summary, to the best of our knowledge, MAGA-GLTrans represents the first approach in computational pathology that simultaneously addresses classification capability, computational efficiency, storage requirements, and communication burden. By significantly enhancing efficiency, MAGA-GLTrans renders the computational process virtually imperceptible to pathologists, marking a critical step toward integrating computational pathology into real-world clinical practice. This advancement is particularly impactful in applications such as computer-assisted telepathology (CATP) for intraoperative frozen section diagnosis, where processing speed, accuracy, and resource efficiency are paramount.

### v2

@@ Results
Figure 4: Extensibility of MAG to the histopathology-specific encoder. MAG can be applied to a histopathology-specific encoder or pathology vision foundation model to extract fine-grained domain-specific features from low magnification inputs. Receiver operating characteristic (ROC) curves demonstrate the performance of CLAM (a, d, g), TransMIL (b, e, h) and MAG-GLTrans (c, f, i) on TCGA-NSCLC using MAG trained with domain-unspecific (ImageNet pre-trained ResNet50) and domain-specific models (TCGA pre-trained CTransPath), respectively. From top to bottom: ROC curves at 20×20$\times$20 ×, 5×5$\times$5 × without MAG and 5×5$\times$5 × with MAG.
MAG can be extended to the histopathology-specific encoder
@@
Histopathology-specific encoders or foundation models [24, 23] have been proven to be much more effective for WSI analysis in various CPath tasks compared with the feature extractor pre-trained on natural images. MAG, as a self-supervised model that can align the learned knowledge and restore the information loss between two magnifications, is capable of being extended to the histopathology-specific encoder, making it possible for low magnification inputs. Here, we investigate the extensibility of MAG on the earliest histopathology-specific model, CTransPath [24] as well as the recently proposed pathology vision foundation model, UNI [23]. We conduct this experiment on the NSCLC subtyping task (TCGA-NSCLC). CLAM, TransMIL and MAG-GLTrans are trained at the high magnification 20×20$\times$20 × and the low magnification 5×5$\times$5 × (with and without MAG).
@@
As we expected, the histopathology-specific pre-trained feature extractors at 20×20$\times$20 × brings appropriate improvements for all three models (Figure 4a-c) compared with the ImageNet pre-trained feature extractor with respect to AUC (CTransPath: CLAM=0.9631, $+7.01\%$, TransMIL=0.9536, $+3.52\%$, MAG-GLTrans=0.9659, $+0.83\%$) and (UNI: CLAM=0.9790, $+8.06\%$, TransMIL=0.9690, $+5.06\%$, MAG-GLTrans=0.9673, $+0.97\%$). When reducing the magnification level to 5×5$\times$5 × without the support of MAG, the classification performance (Figure 4d-f) with three different encoders of the three WSI analysis models decreases due to the information loss. Even though CTransPath without MAG also faces an obvious performance degradation at 5×5$\times$5 ×, applying a histopathology-specific feature extractor is still a better solution compared with the ImageNet pre-trained one. When training MAG on three different encoders at the 5×5$\times$5 × magnification level, the performance of three WSI analysis architectures returns to a favorable level, even outperform the model trained at 20×20$\times$20 × with the ImageNet pre-trained feature extractor. The model-agnostic nature of MAG is a highly significant feature that enables it to be seamlessly integrated with any top-tier, histopathology-specific pre-trained model, resulting in minimal information loss. This capability substantially enhances computational efficiency and communication efficiency, all without compromising the model’s classification performance.
@@
Figure 5: Visualization of the features extracted by MAG (TCGA-NSCLC). a, The prediction results (top) of one fold on the subtyping task (TCGA-NSCLC, $n=106$) under different magnification levels. Patients are sorted by the ground truth labels. Heatmaps (bottom) demonstrate the similarity of the features between high magnification level and low magnification level (with and without MAG). We show the feature similarity in slide level, which is the mean of all the patches. Red represents higher feature similarity. b, A sample case WSI (left) of LUSC with pathologist-annotated carcinoma region (yellow border), visualization of the unsupervised clustering of the patch-level features (top right) and the heatmap (bottom right) at 20×20$\times$20 ×, 5×5$\times$5 × with MAG and 5×5$\times$5 × without MAG. c, T-SNE visualization of the image embeddings from three different models in the Kather colon dataset [39] with nine tissue types.
MAG features are reliable and interpretable
@@
Model interpretability and reliability are the most important natures when applying AI models to real-world clinical scenarios. As a feature extractor given low magnification inputs, MAG faces a significant challenge of information loss, potentially undermining the model’s reliability and interpretability. In this experiment, we investigate the feature quality with and without MAG, by visualizing them at cohort, slide and patch levels. The visualization of NSCLC subtyping (TCGA-NSCLC) and breast cancer lymph-node metastasis detection (CAMELYON16) are shown in Figure 5 and Extended Figure 4 respectively.
@@
To evaluate and visualize whether MAG is able to restore the information loss at the low magnification, we quantify the similarity of the features between low magnification 5×5$\times$5 × (with and without MAG) and high magnification 20×20$\times$20 ×. Specifically, we calculate a slide-level similarity score by averaging the similarity scores (normalized Euclidean distances) across all patches. The heatmaps depicted in Figure 5a indicate that the features extracted from the low magnification images using MAG are significantly more similar to those from the high magnification images than the features extracted at low magnification without MAG. This finding suggests that MAG can refine the features extracted from low magnification images, effectively bridging a substantial information gap of up to 16 times. With those more reliable features, the rate of misclassify has noticeably decreased at 5×5$\times$5 × magnification using MAG compared with the ones in the 5×5$\times$5 × without MAG. We also observe that the misclassified samples at 5×5$\times$5 × without MAG tend to exhibit lower feature similarities, typically falling below the threshold of 0.5. For the CAMELYON16 dataset with fewer training samples and larger color variations, unreliable features can severely impair the prediction model, resulting in significantly degraded classification performance with more misclassified samples (Extended Figure 4).
@@
To further visualize the robustness and the representativeness of MAG features, and to evaluate whether those features deliver enough semantic information, we present a slide-level visualization in Figure 5b. We first apply an unsupervised clustering approach (K-means) on the patch-level features and group them into four clusters. The cluster maps shown in the top-right row demonstrate that the features extracted from high magnification patches are capable of capturing and reflecting the morphological characteristics accurately with a clear boundary between tumor and normal tissues and a high agreement with the boundary drawn by the pathologist. When reducing the magnification to 5×5$\times$5 ×, the tissue boundaries become unclear and the clustering map is much noisier than the high magnification one. This observation indicates that reducing image resolution can greatly harm the semantic information, resulting in unstable prediction results. By bridging information gap using MAG, the clustering map of 5×5$\times$5 × with MAG maintains a high visual consistency with the one of 20×20$\times$20 ×. Next, we plot the heatmaps of the WSI analysis model GLTrans in the top-right row in Figure 5b. The color intensity of the heatmaps correlates with the confidence of the classification outcome for that particular patch. When without MAG, the lower magnification features lead to less effective heatmaps, indicating that the feature representation of the predictive models is impaired by these unstable and less detailed features. Applying MAG significantly enhances the effectiveness of features, resulting in heatmaps that are much more consistent with those at higher magnification. The feature space of the image embeddings (Figure 5c) also demonstrates the effectiveness of MAG.
@@
Figure 6: MAG-GLTrans on telepathology for intraoperative frozen section diagnosis. a, Local hospital transfers low magnification WSIs to the cloud server and obtains a diagnostic result and an interpretable heatmap in a only few seconds. b-f, Statistical results (b) and the box plots at different magnification levels with respect to the pixel quantity (c), data transfer time (d), model inference time (e) and the diagnostic performance of AUC (f). h, A sample case and its corresponding heatmap as well as the patches with high confidence (red border) and low confidence (blue border).
@@
We evaluate the potential of MAG-GLTrans for computer-assisted telepathology in intraoperative frozen section diagnosis, a critical need for hospitals lacking on-site expert pathologists, with a focus on classification performance, computational efficiency, and communication efficiency. This investigation is conducted on the NSCLC subtyping task using the TCGA-NSCLC dataset (frozen sections), which includes 1067 frozen section WSIs of LUAD and 1100 WSIs of LUSC. A 10-fold cross-validation approach is employed, with all patients randomly partitioned into a training set (80%), a validation set (10%), and a test set (10%). To prevent data leakage, all slides from a single patient are assigned to the same set. Classification performance is assessed using the AUC metric. For each WSI, we store each magnification level as a separate file to accurately measure storage consumption, data transfer time, and model inference time across different magnifications. Transmission performance is evaluated under a 200 Mbps bandwidth connection during typical working hours (8:00-18:00) to simulate real-world clinical conditions.
@@
For local hospitals lacking on-site expert pathologists, particularly in underdeveloped regions or countries, we propose an efficient computer-assisted telepathology (CATP) framework tailored for frozen section diagnosis (Figure 6a). This framework enables local hospitals to transmit low-magnification whole slide images (WSIs) to cloud servers and receive computer-assisted diagnostic results alongside interpretable heatmaps within seconds. To rigorously assess the framework’s efficacy, we conducted an ablation study, as depicted in Figure 6b. The results reveal that, as anticipated, higher magnification WSIs require significantly more storage due to their substantially larger pixel counts (20$\times$: 0.503 Gigapixels/WSI vs. 5$\times$: 0.033 Gigapixels/WSI), leading to increased data transfer and model inference times. On average, processing a single WSI takes approximately one minute (59.13 seconds) for file transmission and half a minute (32.32 seconds) for model inference. In real-world clinical settings, these times can vary considerably due to factors such as network bandwidth, latency fluctuations, computational resource availability, concurrent user load, and other environmental variables. Reducing magnification yields remarkable efficiency gains, improving processing speed by more than tenfold. Specifically, data transfer time drops dramatically from 59.13 seconds per WSI at 20$\times$×magnification to just 2.79 seconds at 5$\times$, while model inference time decreases from 32.32 seconds to around 5 seconds per WSI. However, this optimization comes at a significant cost to diagnostic accuracy, as indicated by the sharp decline in AUC values from 0.8995 at 20$\times$ to 0.6267 at 5$\times$, severely limiting clinical utility. Our proposed MAG-GLTrans framework effectively addresses this trade-off by maintaining high predictive performance (AUC: 0.8418) with only marginal degradation compared to 20$\times$ magnification, while simultaneously mitigating the drawbacks of magnification reduction. By significantly reducing both data transfer and model inference times, this innovative CATP framework greatly enhances the clinical feasibility of intraoperative frozen section diagnosis.
@@ Discussion
We introduce MAG-GLTrans, a general-purpose WSI analysis architecture, designed for a wide range of computational pathology applications, including cancer subtyping, metastasis detection, and mutation prediction. Beyond its exceptional classification capabilities, MAG-GLTrans stands out for its computational efficiency, storage-friendly design, extensibility, and communication efficiency. These attributes, whose importance is first emphasized in this study, are often overlooked in the field of computational pathology (CPath). The primary obstacle to advancing CPath towards greater efficiency lies in the fact that current WSI analysis architectures typically process WSIs at high magnifications [40]. This approach is constrained by the limitations of existing computer hardware. While directly reducing image resolution can improve efficiency, it often comes at the cost of performance due to information loss. MAG addresses this challenge by bridging the information gap and extracting fine-grained features from low-magnification images, thereby achieving an optimal balance between performance and efficiency. Through 10-fold cross-validation experiments on five CPath tasks, our model demonstrates robust classification performance at 5$\times$×magnification, comparable to two state-of-the-art WSI analysis architectures operating at 20$\times$, while reducing computational time by up to 10.7 times, storage overhead by up to 28.6 times, and communication time by up to 21.5 times. We believe this represents a significant and milestone finding that brings CPath technology one step closer to real-world clinical adoption. For instance, a recent study highlights that reviewing WSIs results in a median overall 19% decrease in efficiency per case compared to glass slides [41]. This inefficiency is further exacerbated when pathologists must wait tens of seconds for a decision on a single slide from CPath systems. Such delays are particularly pronounced when analyzing multiple hematoxylin and eosin (H&E) stained diagnostic slides and immunohistochemistry slides, a common scenario for patients with larger tumors. Even if WSIs are processed offline before pathologist review, hospitals with high patient volumes would require substantial computational resources. MAG-GLTrans addresses this issue by reducing computation turnaround time to just a few seconds, making the computational process virtually seamless for pathologists.
@@
In addition to enhancing local pathology diagnosis, we also investigate the potential of MAG-GLTrans in enabling telepathology (TP), a field of pathology that utilizes telecommunication technologies to facilitate the transfer of digital pathology slides across geographically distant locations [42]. Intraoperative frozen section diagnosis stands out as one of the most promising applications of TP, particularly in settings where pathologists are scarce, such as hospitals without on-site pathology expertise [43, 44, 45]. Integrating computational pathology techniques to enable computer-assisted telepathology (CATP) for intraoperative diagnosis has the potential to improve diagnostic accuracy [46]. However, the unpredictable data transfer times under complex network conditions and the prolonged inference times associated with processing gigapixel WSIs pose significant challenges, potentially increasing turnaround times and compromising diagnostic efficiency [25]. These issues highlight the need for further validation of CATP’s feasibility. Our evaluation on the frozen section diagnosis dataset (TCGA-NSCLC (frozen sections)) demonstrates that MAG-GLTrans can significantly reduce the turnaround time required for both data transfer and computation.
@@
Due to the absence of domain-specific encoders before the emergence of foundation models, earlier WSI analysis pipelines typically relied on ImageNet pre-trained models to extract patch-level features as a compromise solution. However, with the accumulation of large-scale histopathology datasets and advancements in computational resources, self-supervised learning on histopathology datasets [24] and the development of visual foundation models [13, 23, 47] have led to significant progress in computational pathology. Notably, histopathology-specific encoders such as CTransPath [24], UNI [23], PLIP [13], Prov-GigaPath [11], MUSK [48] and Virchow [49] have demonstrated superior effectiveness across various CPath tasks compared to encoders trained on natural images. To preserve fine-grained information, most existing histopathology-specific encoders are still trained on high-magnification images. MAG, however, can be extensible to any histopathology-specific encoder in a self-supervised manner, enabling the extraction of fine-grained features even from low-magnification inputs. Experimental results indicate that MAG, when trained with a histopathology-specific encoder, further enhances classification performance compared to MAG trained with ImageNet. Additionally, due to its plug-and-play nature, MAG can be seamlessly integrated into other WSI analysis architectures, inheriting their strengths while improving their efficiency. For example, when MAG is applied as the feature extractor for CLAM, the performance remain largely unaffected even at low magnification. This exceptional extensibility of MAG significantly extends its lifecycle, particularly in light of the rapid pace of algorithmic advancements and iterations in the field.
@@
While the MAG-GLTrans model has demonstrated remarkable efficiency and accuracy in handling low-magnification inputs, several areas warrant further exploration to unlock its full potential. Firstly, we hypothesize that the model’s capabilities could extend beyond its current resolution limits. By applying the model to even lower magnifications, such as 2.5$\times$, it may be possible to achieve further improvements in computational speed and communication efficiency. However, this approach requires a careful balance between reducing image resolution and preserving essential diagnostic details. Whether this trade-off will compromise the model’s effectiveness at lower magnifications remains an open question that merits thorough investigation. Additionally, due to the lack of large-scale datasets, MAG was trained on a limited number of WSIs as a compromise. Investigating its performance on a larger-scale dataset would be highly valuable, as it could further enhance the model’s generalization capabilities and robustness. This exploration could provide deeper insights into the scalability and effectiveness of MAG when applied to more diverse and extensive histopathology data, potentially unlocking even greater performance improvements and broader applicability in computational pathology.
@@
In summary, to the best of our knowledge, MAG-GLTrans represents the first approach in computational pathology that simultaneously addresses classification capability, computational efficiency, storage requirements, and communication burden. By significantly enhancing efficiency, MAG-GLTrans renders the computational process virtually imperceptible to pathologists, marking a critical step toward integrating computational pathology into real-world clinical practice. This advancement is particularly impactful in applications such as computer-assisted telepathology (CATP) for intraoperative frozen section diagnosis, where processing speed, accuracy, and resource efficiency are paramount.